# Decompose-and-Compose: A Compositional Approach to Mitigating Spurious Correlation

### v1

@@ Abstract
While standard Empirical Risk Minimization (ERM) training is proven effective for image classification on in-distribution data, it fails to perform well on out-of-distribution samples. One of the main sources of distribution shift for image classification is the compositional nature of images. Specifically, in addition to the main object or component(s) determining the label, some other image components usually exist, which may lead to the shift of input distribution between train and test environments. More importantly, these components may have spurious correlations with the label. To address this issue, we propose Decompose-and-Compose (DaC), which improves robustness to correlation shift by a compositional approach based on combining elements of images. Based on our observations, models trained with ERM usually highly attend to either the causal components or the components having a high spurious correlation with the label (especially in datapoints on which models have a high confidence). In fact, according to the amount of spurious correlation and the easiness of classification based on the causal or non-causal components, the model usually attends to one of these more (on samples with high confidence). Following this, we first try to identify the causal components of images using class activation maps of models trained with ERM. Afterward, we intervene on images by combining them and retraining the model on the augmented data, including the counterfactual ones. Along with its high interpretability, this work proposes a group-balancing method by intervening on images without requiring group labels or information regarding the spurious features during training. The method has an overall better worst group accuracy compared to previous methods with the same amount of supervision on the group labels in correlation shift.
@@ 1 Introduction
0-0.250.25-0.50.5-0.750.75-1000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111quantiles of loss(a)$C$ $S$ Score00.51Refer to caption $C=0.12$ $S=0.88$ Refer to caption $C=0.12$ $S=0.88$ Refer to caption $C=0.88$ $S=0.12$ Refer to caption $C=0.88$ $S=0.12$ pt(b)0-0.250.25-0.50.5-0.750.75-1000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111quantiles of loss(c)$C$ $S$ Score00.51Refer to caption $C=0.79$ $S=0.30$ Refer to caption $C=0.79$ $S=0.32$ Refer to caption $C=0.55$ $S=0.32$ Refer to caption $C=0.32$ $S=0.28$ pt(d)
Figure 1: Behaviour of a model trained with standard ERM in different datasets. Based on the easiness of inferring the label from the causal or non-causal parts across the whole dataset, the model attends more to one of them, this behaviour is more evident on samples on which the model has a low loss. (a), (b) Average xGradCAM score of Cifar10 (causal) and and MNIST (non-causal) pixels in four loss quantiles of the Dominos train set. The model generally attends more to the non-causal parts, and as the loss decreases, the non-causal attention increases. (c), (d) Average xGradCAM score of foreground (causal) and background (non-causal) pixels in four loss quantiles of the Waterbirds train set. The model generally pays attention to the causal parts, and as the loss decreases, the causal attention increases.
@@
Another line of work, approaches this problem by balancing minority/majority groups of data [7, 5, 23] to remove spurious correlation. Among these works, [7] proposes DFR, which retrains the last layer of a model previously trained by ERM, with group-balanced data to make it robust to spurious correlation. Nonetheless, DFR requires group labels. On the other hand, methods like [11, 18, 15] try to learn a robust model by reweighting samples that are less likely to contain spurious attributes, without access to group labels during training. Most of these methods rely on the assumption that datapoints on which the model has a high loss are most probably from minority groups and want to place more emphasis on these samples. However, in these samples, the obscure core object may be the source of high loss (i.e., the target object itself cannot be easily classified), and overrating these samples may have some side effects. To be more precise, we need to discover and analyze parts of images to make a more accurate decision.
Due to the compositional nature of images, the problem of correlation shift can be viewed through the lens of compositionality, as models fall into the trap of spurious correlation because they make their predictions based on non-causal components of images. This fine-grained perspective could lead to a more precise approach compared to methods that consider images as a whole. While the viewpoint of compositionality is essential to OOD generalization, especially when facing correlation shift, only a limited number of works have explored this problem from this perspective. As a recent work, [34] combines different images and uses them for model distillation on the representation level. Nonetheless, they cannot label the combined images, as they could not determine whether the parts taken from images for combining are causal or non-causal parts. Additionally, they do not offer any evaluations on correlation shift benchmarks. Masktune [3] takes a step further and hypothesizes that in datasets exhibiting spurious correlation, the parts of an image with high attribution scores are non-causal and misleading, and based on this assumption, masks these parts for finetuning a model.
@@
Figure 2: (a) Image as a composition of causal and non-causal components. (b) The edge between $S$ and $\tilde{S}$ can be removed by intervention on components of $\tilde{S}$. This removes the spurious correlation between $\tilde{S}$ and $Y$.
In this paper, we first analyze the behavior of a model trained with ERM and utilize its attribution map on images to decompose them into causal and non-causal parts; then, based on the performance of the model trained with our method on the validation set, we identify the causal parts in images.
More precisely, as opposed to MaskTune [3], which assumes that for a given model trained with standard ERM, the regions with high attribution scores are spurious ones, we show that most of the attention regions of a model trained with ERM are usually either on causal or on non-causal parts of images, based on the easiness of predicting the label from the causal and non-causal parts respectively. Gaining this knowledge about the causal parts enables us to intervene on datapoints from the majority groups to create new ones from under-represented groups as a means of group balancing.
@@
  • We provide a thorough analysis of the behavior of models trained with standard ERM, especially on low loss data.
@@ 2 Preliminaries
Spurious correlations that are perceivable by humans can be categorized into two groups: 1) Spurious parts: spurious correlations between other objects of the image and the label (e.g., spurious correlation of the background and the label in the Waterbirds [22] dataset) and 2) Spurious attributes: spurious correlations between some non-causal attributes of the object of interest and the label (e.g., spurious correlation between the color of the digits and their label in the CMNIST [2] dataset). In this section, we propose a method that provides robustness to the first type of spurious correlation, which is prevalent in most benchmarks. We propose a compositional view to mitigate correlation shifts by discovering non-causal parts of images and intervening on these parts.
@@ 2.1 A Causality Viewpoint to Spurious Correlation
We consider a dataset $D=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}$ that is associated with spurious correlations. Our objective is to train a classifier $f$ that achieves consistent performance across both the training distribution and test distributions containing shifts. We assume that the primary difference between distributions is the shift in the spurious correlations.
To study this problem from the perspective of causality, we model the data-generating process as the SCM [17] illustrated in Fig. 1(a). In this SCM, $C$ and $S$ indicate unobservable causal and non-causal variables, from which the observable causal and non-causal components $\tilde{C}$ and $\tilde{S}$ for an image are obtained. The final image $X$ is the output of $\psi(\tilde{C},\tilde{S})$, where $\psi(.,.)$ is a combining function. The label of the image is caused by $\tilde{C}$.
In the case of typical distribution shift studied in the literature, a hidden confounder $E$, mostly referred to as the environment [2] or group [22] variable in the literature, would be present in the SCM such that $S\leftarrow E\rightarrow C$. This creates the backdoor path $\tilde{S}\leftarrow S\leftarrow E\rightarrow C\rightarrow\tilde{C}\rightarrow{Y}$, which introduces a spurious correlation between $\tilde{S}$ and $Y$. In the case of spurious correlation, $E$ is mainly sample selection bias.
@@
Another solution that is effective even in the absence of group annotation is removing the edge $S\xrightarrow{}\tilde{S}$ by intervening on some components in $\tilde{S}$ in order to break the backdoor path, as shown in Fig. 1(b). More concretely, this solution intervenes on a subset of non-causal components of images without changing the label in order to reduce their correlation. Intervention on $\tilde{S}$ could be done in a more efficient manner if we could set $\tilde{S}$ to a value that has less co-occurrence with $Y$. Such intervention would create a new datapoint that can be assigned to a minority group. Hence, this type of intervention would be a method for upweighing datapoint from minority groups.
@@ 3 To Which Does ERM Attend More?
To intervene on the non-causal components of an image, it is essential to determine the causal and non-causal parts of it first. The attribution map of a model trained with standard ERM on an image could be utilized for distinguishing these parts. MaskTune [3] partly addressed this issue by assuming that for a model trained with standard ERM on a dataset exhibiting spurious correlations, the image parts with high attribution scores are non-causal. However, this assertion does not hold for the majority of realistic datasets. Specifically, the behavior of a model trained with ERM might vary depending on the easiness of predicting labels from the causal and non-causal parts across different datasets.
Since non-causal parts of images that have spurious correlations with the label, such as the background in the Waterbirds dataset, are shortcuts for models, it is often presumed that a model trained with standard ERM attends more to the non-causal parts of images. However, we show that this assumption doesn’t hold in many real-world scenarios. Across the entire dataset, as opposed to the causal components, non-causal ones do not persistently appear in accordance with the label. Consequently, the causal parts may become generally more predictive. As a result, the easiness of predicting labels from the causal and non-causal parts across the entire dataset determines the focus of ERM.
To illustrate this point, we report the average xGradCAM [25] score for the foreground and background pixels in the Waterbirds datasets in Fig. 1(a,b) and for Cifar part and MNIST part of Dominoes dataset [16] in Fig. 1(c,d). For the Waterbirds, the average score of the foreground (causal) pixels exceeds that of the background (spurious part) and for the Dominoes, the average score of the spurious part (MNIST part) exceeds that of the causal pixels (i.e. Cifar part). Indeed, if the spurious patterns are easy to learn, the model may pay more attention to the spurious parts. This tendency becomes more pronounced in samples with low loss, as depicted in Fig. 1.
@@ 4 Method
Regardless of whether models trained with ERM attend more to the causal or non-causal parts of an image, their attribution map on the image remains useful for distinguishing these two parts. In the following, we introduce a method for identifying the most predictive parts of images using these attribution maps.
@@ 4.1 Adaptive Masking by ERM
Given a model $f_{\theta}(.)$ that is trained with ERM, and a datapoint $(x,y)$ on which $f$’s loss is low, we want to find a mask that conceals pixels in $x$ except for the most predictive ones. The assumption that the loss on $(x,y)$ is low is necessary, since we are more confident that the predictive parts in $x$ are indeed predictive of the correct label.
@@
Figure 3: Adaptive masking according to the attention scores obtained from the ERM model. The loss value of the masked images for which different portion $p$ of pixels (with the lowest attention score) has been masked is shown as $l_{p}$. (a) The loss curve for an image of the Dominoes dataset with the label ’truck’ on which ERM has non-causal attention, and (b) The loss curve for an image of the MetaShift dataset with label ’dog’ on which ERM has causal attention.
@@
Figure 4: (a) An overview of our DaC method. For each batch, a $q$ portion of samples with the lowest loss are selected. Then images of different labels are combined together by the Mask and Combine module. The overall loss to update the model’s last layer parameters is the sum of the loss on the original batch ($L_{CE}$) and the combined data ($L_{\text{comb}}$). The algorithm for this method is shown in Algorithm 2. (b) The Mask and Combine module. The two input images $x^{(i)}$ and $x^{(j)}$ are masked by Algorithm 1. Afterwards, The selected part of $x^{(i)}$ and the masked parts of $x^{(j)}$ are combined, and the remaining gaps are filled with the mean value of the batch. The new combined image has the same label as $x^{(i)}$, and is used for training the last layer of the model.
@@ 4.2 Decompose-and-Compose (DaC)
Viewing images as a combination of components, gives us the upperhand of being able to intervene [17] on images by combining components of different images in order to create novel combinations less seen by the model. To be more specific, in the case of spurious correlation, by combining components of a number of selected inputs, novel images that are more similar to underrepresented data could be created and used during the training of the model as a means of upweighing the underrepresented group. Here we propose a method for obtaining new datapoints from minority groups, by combining the ones from the majority groups. Given dataset $\mathcal{D}$, we consider a pretrained classifier $f_{\theta}(x)=w\circ g_{\phi}(x)$ with standard ERM on $\mathcal{D}$, in which $g_{\phi}$ is a feature extractor and $w$ is a linear predictor. Afterward, inspired by [7], we intend to retrain only $w$ on an augmented variation of $\mathcal{D}$ that is prepared by intervening on samples to produce new ones, in order to upweigh the underrepresented or minority group.
Selecting low loss examples and decomposing them. For a given training batch $\mathcal{B}$, we first select subset =$\mathcal{B}^{\prime}=\{(x^{(i)},y^{(i)})\in\mathcal{B}|(x^{(i)},y^{(i)})$ is among $q$ portion of samples with the lowest loss} and $q$ is the hyperparameter denoting the porion of the selected samples. For each sample $(x^{(i)},y^{(i)})\in\mathcal{B}^{\prime}$, first, by Algorithm 1, the mask $m(x^{(i)})$ will be obtained. Then, two parts of image $x$ is found as $x^{(i)}\odot m(x^{(i)})$ and $x^{(i)}\odot(1-m(x^{(i)}))$. But still, we do not know which of these two parts consists more of the causal regions.
As explained in Sec. 3, based on the predictive power of the causal and non-causal parts of the images, the model may pay more attention to one of them . We inspect both of these assumptions and find which of them yields better results on the validation data. Therefore, we consider the $causalflag$ as a hyperparameter in Algorithm 2, and when this flag indicates the non-causal assumption, masks are reversed as shown in Lines 10-13 to obtain more causal regions.
Composing the causal and non-causal parts of two images. Given a sample $(x^{(i)},y^{(i)})\in\mathcal{B}^{\prime}$, another sample $(x^{(j)},y^{(j)})\in\mathcal{B}^{\prime}$ is selected randomly such that $y^{(i)}\neq y^{(j)}$. Then, a combination of the two images is computed as below:
@@
Where $b$ is a $1\times 1\times 3$ vector indicating the mean of $\mathcal{B}$ across the color channels. This formula constructs the combined image by putting the selected parts of $x^{(i)}$ and masked parts of $x^{(j)}$ that is not located on the selected parts of $x^{(i)}$ together, and filling the remaining parts of the image by the default background $b$. The reason for setting the remaining areas equal to $b$ to retain the statistics of the batch as much as possible. Finally, we define =$\mathcal{M}=\{(\hat{x}^{(i)}_{\text{comb}_{j}},y^{(i)})|(x^{(i)},y^{(i)})\in\mathcal{B}^{\prime}\}$ as the combined samples.
As it will be shown in the Appendix, most low-loss datatpoints are the ones from the majority groups. By combining the causal and non-causal parts of two majority datapoints from different labels, we make new datapoints from minority groups, thus group-balance the training data without access to group annotation. More precisely, suppose that the model attends more to the causal parts accross the dataset. Consider $(x^{(i)},y^{(i)})$ and $(x^{(j)},y^{(j)})$ as two samples from majority groups where $x^{(i)}=\psi(\tilde{c}^{(i)},\tilde{s}^{(i)})$ and $x^{(j)}=\psi(\tilde{c}^{(j)},\tilde{s}^{(j)})$ and $y^{(i)}\neq y^{(j)}$. Therefore, $\tilde{s}^{(i)}$ and $\tilde{s}^{(j)}$ are in accordance with $y^{(i)}$ and $y^{(j)}$ respectively. Now, in the combined datapoint $\hat{x}^{(i)}_{\text{comb}_{j}}=\psi(\tilde{c}^{(i)},\tilde{s}^{(j)})$, the non-causal part $\tilde{s}^{(j)}$ does not have a spurious value corresponding to $y^{(i)}$, which makes this datapoint from the minority groups. It is worth mentioning that this combination tends to be a minority sample independent of whether $m(x^{(i)})\odot x^{(i)}$ or $(1-m(x^{(i)}))\odot x^{(i)}$ reflects the cause more.
@@
$L_{\text{total}}=L_{CE}+\alpha L_{\text{comb}}.$ (4)
in which $l$ is the cross-entropy loss and $\alpha$ is the hyperparameter determining the effectiveness of combination. An overview of our Decompose and Compose (DaC) method in addition to its algorithm are shown in Fig. 4 and Algorithm 2 respectively.
Input: Model $f_{\theta}(.)$; Dataset $\mathcal{D}$; Loss function $l(.,.)$; Hyperparameters, $\alpha,q$, causalflag
1 for epoch=$1,2,\dots K$ do
2       for batch $\mathcal{B}$ in $\mathcal{D}$ do

### v3

@@ Abstract
While standard Empirical Risk Minimization (ERM) training is proven effective for image classification on in-distribution data, it fails to perform well on out-of-distribution samples. One of the main sources of distribution shift for image classification is the compositional nature of images. Specifically, in addition to the main object or component(s) determining the label, some other image components usually exist, which may lead to the shift of input distribution between train and test environments. More importantly, these components may have spurious correlations with the label. To address this issue, we propose Decompose-and-Compose (DaC), which improves robustness to correlation shift by a compositional approach based on combining elements of images. Based on our observations, models trained with ERM usually highly attend to either the causal components or the components having a high spurious correlation with the label (especially in datapoints on which models have a high confidence). In fact, according to the amount of spurious correlation and the easiness of classification based on the causal or non-causal components, the model usually attends to one of these more (on samples with high confidence). Following this, we first try to identify the causal components of images using class activation maps of models trained with ERM. Afterwards, we intervene on images by combining them and retraining the model on the augmented data, including the counterfactual ones. This work proposes a group-balancing method by intervening on images without requiring group labels or information regarding the spurious features during training. The method has an overall better worst group accuracy compared to previous methods with the same amount of supervision on the group labels in correlation shift. Our code is available at https://github.com/fhn98/DaC.
@@ 1 Introduction
0-0.250.25-0.50.5-0.750.75-1000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Quantiles of Loss(a)$C$ $S$ Score00.51Refer to caption $C=0.12$ $S=0.88$ Refer to caption $C=0.12$ $S=0.88$ Refer to caption $C=0.88$ $S=0.12$ Refer to caption $C=0.88$ $S=0.12$ (b)0-0.250.25-0.50.5-0.750.75-1000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Quantiles of Loss(c)$C$ $S$ Score00.51Refer to caption $C=0.79$ $S=0.30$ Refer to caption $C=0.79$ $S=0.32$ Refer to caption $C=0.55$ $S=0.32$ Refer to caption $C=0.32$ $S=0.28$ (d)
Figure 1: Behaviour of a model trained with standard ERM in different datasets. Based on the easiness of inferring the label from the causal or non-causal parts across the whole dataset, the model attends more to one of them, this behaviour is more evident in samples on which the model has a low loss. (a), (b) Average xGradCAM score of Cifar10 (causal) and MNIST (non-causal) pixels in four loss quantiles of the Dominoes training set. The model generally attends more to the non-causal parts, and as the loss decreases, the non-causal attention increases. (c), (d) Average xGradCAM score of foreground (causal) and background (non-causal) pixels in four loss quantiles of the Waterbirds training set. The model generally attends to the causal parts, and as the loss decreases, the causal attention increases.
@@
Another line of work approaches this problem by balancing minority/majority groups of data [8, 6, 25] to remove spurious correlation. Among these works, [8] proposes DFR, which retrains the last layer of a model previously trained by ERM, with group-balanced data to make it robust to spurious correlation. Nonetheless, DFR requires group labels. On the other hand, methods like [13, 20, 17] try to learn a robust model by upweighting samples that are less likely to contain spurious attributes, without access to group labels during training. Relying on the assumption that datapoints on which the model has a high loss are most probably from minority groups, most of these methods aim to place more emphasis on these samples. However, in these samples, the obscure core object may be the source of high loss (i.e., the target object itself cannot be easily classified), and overrating these samples may have some side effects. To be more precise, we need to discover and analyze parts of images to make a more accurate decision.
Due to the compositional nature of images, the problem of correlation shift can be viewed through the lens of compositionality, as models fall into the trap of spurious correlation because they make their predictions based on non-causal components of images. This fine-grained perspective could lead to a more precise approach compared to methods that consider images as a whole. While the viewpoint of compositionality is essential to OOD generalization, especially when facing correlation shift, only a limited number of works have explored this problem from this perspective. As a recent work, [36] combines parts of different images and uses them for model distillation on the representation level. Nonetheless, they cannot label the combined images, as they could not determine whether the parts taken from images for combining are causal or non-causal parts. Additionally, they do not offer any evaluations on correlation shift benchmarks. Masktune [3] takes a step further and hypothesizes that in datasets exhibiting spurious correlation, the parts of an image with high attribution scores according to a model trained with ERM, are non-causal and misleading, and based on this assumption, masks these parts for finetuning the model.
@@
Figure 2: (a) Image as a composition of causal and non-causal components. (b) The edge between $S$ and $\tilde{S}$ can be removed by intervention on components in $\tilde{S}$. This removes the spurious correlation between $\tilde{S}$ and $Y$.
In this paper, we first analyze the behaviour of a model trained with ERM and utilize its attribution map on images to decompose them into causal and non-causal parts; then, based on the performance of the model trained with our method, on the validation set, we identify the causal parts.
More precisely, as opposed to MaskTune [3], which assumes that for a given model trained with standard ERM, the regions with high attribution scores are spurious ones, we show that a model trained with ERM usually focuses on either causal or non-causal parts of images, based on the easiness of predicting the label from them. Gaining this knowledge about the causal parts enables us to intervene on datapoints from the majority groups to create new ones from under-represented groups as a means of group balancing.
@@
  • We provide an analysis of the behaviour of models trained with standard ERM, especially on low-loss data.
@@ 2 Preliminaries
Spurious correlations that are perceivable by humans can be categorized into two groups: 1) Spurious parts: spurious correlations between other objects of the image and the label (e.g., spurious correlation of the background and the label in the Waterbirds [24] dataset) and 2) Spurious attributes: spurious correlations between some non-causal attributes of the object of interest and the label (e.g., spurious correlation between the colour of the digits and their label in the CMNIST [2] dataset). In this section, we propose a method that provides robustness to the first type of spurious correlation, which is prevalent in most benchmarks. The proposed method mitigates correlation shifts by discovering non-causal parts of images and intervening on these parts.
2.1 Problem Definition
Consider a dataset $D_{tr}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}$ for a classification problem. Each $x\in\mathcal{X}$ has core features $c$, which are the cause of the label, and their correlation with the label is persistent across different environments. It has also spurious features $s$. Based on combinations of core and spurious features, $(c,s)$, training samples can be partitioned into several groups. We consider the case when there is an imbalance between the size of groups in the training set. In this case, the group containing the majority of samples in a class is called the majority group of that class, and the others are called the minority groups. This imbalance induces a spurious correlation between the spurious features and the label, i.e. the value of the spurious features and the label corresponding to a majority group are frequently seen together in the dataset. The proportion of samples in groups could be different in the test set, causing a correlation shift between the training and the test sets. For instance, in the Waterbirds dataset [24], in which the task is to determine whether each image shows a waterbird or a landbird, the core and spurious features are the foreground and the background respectively. Waterbirds consists of four groups: waterbirds on water background, landbirds on land background, waterbirds on land background and landbirds on water background, with the first two being the majority groups.
Our objective is to train a classifier, denoted as $f$, that performs well across both the training and test distributions. This entails ensuring that $f$ exhibits strong performance not only on the majority but also on the minority groups.
@@ 2.2 A Causality Viewpoint to Spurious Correlation
To study the problem of spurious correlation from the perspective of causality, we model the data-generating process as the Structural Causal Model (SCM) [19] shown in Fig. 2(a). In this SCM, $C$ and $S$ indicate unobservable causal and non-causal variables, from which the observable causal and non-causal components $\tilde{C}$ and $\tilde{S}$ for an image are obtained. The final image $X$ is the output of $\psi(\tilde{C},\tilde{S})$, where $\psi(.,.)$ is a combining function. The label of the image is caused by $\tilde{C}$.
In the case of spurious correlation, a hidden confounder $E$, mostly referred to as the environment [2] or group [24] variable in the literature, would be present in the SCM such that $S\leftarrow E\rightarrow C$. This creates the path $\tilde{S}\leftarrow S\leftarrow E\rightarrow C\rightarrow\tilde{C}\rightarrow{Y}$, which introduces a spurious correlation between $\tilde{S}$ and $Y$. $E$ is mainly sample selection bias.
@@
Another solution that is effective even in the absence of group annotation is removing the edge $S\xrightarrow{}\tilde{S}$ by intervening on some components in $\tilde{S}$ in order to break the path, as shown in Fig. 2(b). More concretely, this solution intervenes on a subset of non-causal components of images without changing the label to reduce their correlation. Intervention on $\tilde{S}$ could be done in a more efficient manner if we could set $\tilde{S}$ to a value that has less co-occurrence with $Y$. Such intervention would create a new datapoint that can be assigned to a minority group. Hence, this type of intervention would be a method for upweighting datapoint from minority groups.
@@ 3 To Which Does ERM Attend More?
To intervene on the non-causal components of an image, it is essential to determine the causal and non-causal parts of it first. The attribution map of a model trained with standard ERM on an image could be utilized in distinguishing these parts. MaskTune [3] partly addressed this issue by assuming that for a model trained with standard ERM on a dataset exhibiting spurious correlations, the image parts with high attribution scores are spurious.
However, this assertion does not hold for the majority of realistic datasets. Specifically, the behaviour of a model trained with ERM might vary depending on the easiness of predicting the label from the causal and non-causal parts across different datasets.
Since non-causal parts of images that have spurious correlations with the label, such as the background in the Waterbirds dataset, are shortcuts for models, it is often presumed that a model trained with standard ERM attends more to the non-causal parts of images. However, we show that this assumption does not hold in many real-world scenarios. Across the entire dataset, as opposed to the causal components, non-causal ones do not persistently appear in accordance with the label. Consequently, the causal parts may become generally more predictive. As a result, the easiness of predicting labels from the causal and non-causal parts across the entire dataset determines the focus of ERM.
To illustrate this point, we report the average xGradCAM [27] score for the foreground and background pixels in the Waterbirds datasets in Fig. 1(a,b) and for the Cifar part and the MNIST part of the Dominoes dataset [18] in Fig. 1(c,d). For the Waterbirds, the average score of the foreground (causal) pixels exceeds that of the background (spurious part) and for the Dominoes, the average score of the spurious part (MNIST part) exceeds that of the causal pixels (i.e. Cifar part). Indeed, if the spurious patterns are easy to learn, the model may attend more to the spurious parts. This tendency becomes more pronounced in samples with low loss, as depicted in Fig. 1.
@@ 4 Method
Regardless of whether models trained with ERM attend more to the causal or non-causal parts of an image, their attribution map on the image remains useful for distinguishing these two parts. In the following, we introduce a method for identifying the most predictive parts of images using attribution maps.
@@ 4.1 Adaptive Masking by ERM
Given a model $f_{\theta}(.)$ that is trained with ERM, and a datapoint $(x,y)$ on which $f$’s loss is low, we want to find a mask that conceals pixels in $x$ except for the most predictive ones. The assumption that the loss on $(x,y)$ is low is necessary since we are more confident that the predictive parts in $x$ are indeed predictive of the correct label.
@@
Figure 3: Adaptive masking according to the attention scores obtained from the ERM model. The loss value of the masked images for which different portions $p$ of pixels (with the lowest attention score) has been masked is shown as $l_{p}$. (a) The loss curve for an image of the Dominoes dataset with the label ’truck’ on which the ERM model has non-causal attention, and (b) The loss curve for an image of the MetaShift dataset with the label ’dog’ on which the ERM model has causal attention.
@@
Figure 4: (a) An overview of our DaC method. For each batch, a $q$ portion of samples with the lowest loss is selected. Then images of different labels are combined by the Mask and Combine module. The overall loss to update the model’s last layer parameters is a weighted sum of the loss on the original batch ($L_{CE}$) and the combined data ($L_{\text{comb}}$). The algorithm for this method is shown in Algorithm 2. (b) The Mask and Combine module. The two input images $x^{(i)}$ and $x^{(j)}$ are masked by Algorithm 1. Afterwards, The selected part of $x^{(i)}$ and the masked parts of $x^{(j)}$ are combined, and the remaining gaps are filled with the mean value of the batch. The new combined image has the same label as $x^{(i)}$ and is used for training the last layer of the model.
@@ 4.2 Decompose-and-Compose (DaC)
Viewing images as combinations of components gives us the upper hand of being able to intervene [19] on them. This could be done by combining components of different images to create novel combinations less seen by the model. To be more specific, in the case of spurious correlation, by combining components of different selected inputs, novel images that are more similar to underrepresented data could be created and used during the training of the model as a means of upweighting the underrepresented groups. Here we propose a method for obtaining new datapoints from minority groups, by combining the ones from the majority groups. Given dataset $\mathcal{D}$, we consider a pretrained classifier $f_{\theta}(x)=w\circ g_{\phi}(x)$ with standard ERM on $\mathcal{D}_{tr}$, in which $g_{\phi}$ is a feature extractor and $w$ is a linear predictor. Afterwards, inspired by [8], we intend to retrain only $w$ on an augmented variation of $\mathcal{D}_{tr}$ that is prepared by intervening on samples to produce new ones to upweight the underrepresented or minority groups.
Selecting low-loss examples and decomposing them. For a given training batch $\mathcal{B}$, we first select the subset =$\mathcal{B}^{\prime}=\{(x^{(i)},y^{(i)})\in\mathcal{B}|(x^{(i)},y^{(i)})$ is among $q$ portion of training samples with the lowest loss} and $q$ is the hyperparameter denoting the portion of the selected samples. For each sample $(x^{(i)},y^{(i)})\in\mathcal{B}^{\prime}$, first, by Algorithm 1, the mask $m(x^{(i)})$ will be obtained. Then, the two parts of image $x$ is found as $x^{(i)}\odot m(x^{(i)})$ and $x^{(i)}\odot(1-m(x^{(i)}))$. But still, we do not know which of these two parts consists more of the causal regions.
As explained in Sec. 3, based on the predictive power of the causal and non-causal parts of the images, the model may attend more to one of them . We inspect both of these assumptions and find which of them yields better results on the validation data. Therefore, we consider $causalflag$ as a hyperparameter in Algorithm 2, and when this flag indicates the non-causal assumption, masks are inverted, as shown in Lines 10-13, to obtain more causal regions.
Composing the causal and non-causal parts of two images. Given a sample $(x^{(i)},y^{(i)})\in\mathcal{B}^{\prime}$, another sample $(x^{(j)},y^{(j)})\in\mathcal{B}^{\prime}$ is selected randomly such that $y^{(i)}\neq y^{(j)}$. Then, we combine the two images as below:
@@
where $b$ is a $1\times 1\times 3$ vector indicating the mean of $\mathcal{B}$ across the color channels. This formula constructs the combined image by putting the selected parts of $x^{(i)}$ and masked parts of $x^{(j)}$ that are not located on the selected parts of $x^{(i)}$ together, and filling the remaining parts of the image by the default value $b$. The reason for setting the remaining areas equal to $b$ is to retain the statistics of the batch as much as possible. Finally, we define =$\mathcal{M}=\{(\hat{x}^{(i)}_{\text{comb}_{j}},y^{(i)})|(x^{(i)},y^{(i)})\in\mathcal{B}^{\prime}\}$ as the combined samples.
As will be shown in Appendix 9, most low-loss datatpoints are the ones from the majority groups. By combining the causal and non-causal parts of two majority datapoints from different labels, we make new datapoints from minority groups, thus group-balance the training data without access to group annotation. More precisely, suppose that the model attends more to the causal parts across the dataset. Consider $(x^{(i)},y^{(i)})$ and $(x^{(j)},y^{(j)})$ as two samples from majority groups where $x^{(i)}=\psi(\tilde{c}^{(i)},\tilde{s}^{(i)})$ and $x^{(j)}=\psi(\tilde{c}^{(j)},\tilde{s}^{(j)})$ and $y^{(i)}\neq y^{(j)}$. Therefore, $\tilde{s}^{(i)}$ and $\tilde{s}^{(j)}$ are in accordance with $y^{(i)}$ and $y^{(j)}$ respectively. Now, in the combined datapoint $\hat{x}^{(i)}_{\text{comb}_{j}}=\psi(\tilde{c}^{(i)},\tilde{s}^{(j)})$, the non-causal part $\tilde{s}^{(j)}$ does not have the spurious value corresponding to $y^{(i)}$, which makes this datapoint from the minority groups. It is worth mentioning that this combination tends to be a minority sample independent of whether $m(x^{(i)})\odot x^{(i)}$ or $(1-m(x^{(i)}))\odot x^{(i)}$ reflects the cause more.
@@
$L_{\text{total}}=L_{CE}+\alpha L_{\text{comb}},$ (4)
in which $l$ is the cross-entropy loss and $\alpha$ is the hyperparameter determining the importance of the combinations. An overview of our Decompose and Compose (DaC) method in addition to its algorithm is shown in Fig. 4 and Algorithm 2 respectively.
Input: Model $f_{\theta}(.)=w\circ g_{\phi}(.)$; Dataset $\mathcal{D}_{tr}$; Loss function $l(.,.)$; Hyperparameters $\alpha,q$, causalflag
1
2for epoch=$1,2,\dots K$ do
3       for batch $\mathcal{B}$ in $\mathcal{D}_{tr}$ do